OpenStudy (anonymous):

A dealer paid Php480 for some crates of potatoes. Eight crates spoiled and had to be discarded, but he sold the rest at an increase price of Php3 per crate, gaining Php40. How many crates did he buy?

OpenStudy (dumbcow):

Let x be the number of crates, p be the price paid per crate xp = 480 (x-8)(p+3) = 480+40 ____________________ solve using substitution From 1st equation p = 480/x -> (x-8)((480/x)+3) = 520 distribute 480 +3x -3840/x - 24 = 520 set equal to zero 3x -64 - 3840/x = 0 Multiply equation by x to eliminate x in denominator 3x^2 -64x - 3840 = 0 Use quadratic formula Ignore negative solution x = 48 Therefore, he originally bought 48 crates at 10Php per crate P.S. what is Php? what currency?

OpenStudy (anonymous):

Thank you! thats a lot of help Php is Philippine peso :)

OpenStudy (dumbcow):

ahh, your welcome